# LeanProgress: Guiding Search for Neural Theorem Proving via Proof Progress Prediction

### v1

@@ Abstract
Mathematical reasoning remains a significant challenge for Large Language Models (LLMs) due to hallucinations. When combined with formal proof assistants like Lean, these hallucinations can be eliminated through rigorous verification, making theorem proving reliable. However, even with formal verification, LLMs still struggle with long proofs and complex mathematical formalizations. While Lean with LLMs offers valuable assistance with retrieving lemmas, generating tactics, or even complete proofs, it lacks a crucial capability: providing a sense of proof progress. This limitation particularly impacts the overall development efficiency in large formalization projects. We introduce LeanProgress, a method that predicts the progress in the proof. Training and evaluating our models made on a large corpus of Lean proofs from Lean Workbook Plus and Mathlib4 and how many steps remain to complete it, we employ data preprocessing and balancing techniques to handle the skewed distribution of proof lengths. Our experiments show that LeanProgress achieves an overall prediction accuracy of 75.1% in predicting the amount of progress and, hence, the remaining number of steps. When integrated into a best-first search framework using Reprover, our method shows a 3.8% improvement on Mathlib4 compared to baseline performances of 41.2%, particularly for longer proofs. These results demonstrate how proof progress prediction can enhance both automated and interactive theorem proving, enabling users to make more informed decisions about proof strategies.
Neural Theorem Proving, Automated Reasoning, Lean, LLM Reasoning
@@ 1 Introduction
Formal theorem proving (Avigad, 2023) has emerged as a cornerstone of rigorous mathematical verification, providing machine-checked guarantees for proofs ranging from foundational results (Gowers et al., 2023) to industrial applications (Community, 2022). The Lean proof assistant (Moura & Ullrich, 2021), built on dependent type theory, has witnessed remarkable adoption growth (Best et al., 2023), fueled by collaborative efforts on large-scale formalization projects (mathlib Community, 2020) and novel mathematical developments (Asgeirsson, 2024). This collaborative paradigm shift underscores the urgent need for enhanced tooling to support mathematicians navigating increasingly complex proof environments.
The recent success of Large Language Models (LLMs) in code generation (Rozière et al., 2024) and symbolic reasoning (Yu et al., 2024) has spurred innovations at the intersection of LLMs and formal verification. Some of the works that have been developed include LeanDojo (Yang et al., 2024b) which provides an interactive environment for training LLMs on tactic-level interactions while LLMStep (Welleck & Saha, 2023) and LeanCopilot (Song et al., 2024) focuses on next-tactic suggestion through interface as a useful tool. Lean Agent (Kumarappan et al., 2024) then combines neural suggestion with life-long learning while Lean Aide (Agrawal et al., 2022) and Lean-STaR(Lin et al., 2024) translates statements written in natural language in a doc-string like format to Lean types (including theorem statements) and bootstrapping thoughts. While these systems demonstrate impressive tactic-level accuracy (Johansson, 2023), they primarily optimize for local correctness rather than global proof progress – a critical limitation when navigating Lean’s vast action space (Nawrocki et al., 2023).
@@
Reinforcement learning (RL) presents a theoretically appealing framework for automated theorem proving (Dong et al., 2024), where finding reward signals over proof trajectories is essential. However, the combinatorial explosion of tactic sequences in Lean (Clune, 2023) renders direct RL applications impractical (Setlur et al., 2024). Alphaproof (AlphaProof & AlphaGeometry, 2024) has done RL for theorem proving but it’s not open source and needs enormous compute. Current approaches mitigate this through hybrid architectures (Wang et al., 2023) but remain fundamentally limited by the absence of reliable progress indicators to guide exploration – a prerequisite for effective RL in mathematical domains (Gao et al., 2024).
@@
LeanProgress makes the following key contributions:
Balanced Data Generation Pipeline: We construct a balanced dataset of approximately 80k proof trajectories from Lean Workbook Plus and Mathlib4 by performing a tree search and selecting the shortest path as ground truth. We employ a data balancing strategy based on relative proof progress. Since the useful and non-trivial data of long proofs are long-tailed distributed in the original dataset, we fully utilize long proof data by assigning each state with a remaining step as a label.
Model for Progress Prediction: We fine-tune a DeepSeek Coder V1 1.3b base model to predict the remaining steps, achieving a Mean Absolute Error (MAE) of 3.29 and an overall prediction accuracy of 75.1% on the test set with proof history. Unlike tactic suggestion tools, LeanProgress provides a global view of the proof process by predicting the remaining steps rather than the immediate next tactic.
Progress-Guided Proof Search: We integrate our step prediction model into a best-first search framework. A natural first step for using the progress predictor is combining the predicted remaining steps with the tactic generator’s log probabilities to guide the search. In the future, we hope to use this, instead of just relying on the log probabilities, as a reward for RL. We observe on Mathlib4 a significant improvement of 3.8% with the baseline Reprover performance of 41.2%.
Integration with LeanCopilot: Based on the LeanCopilot framework, we provide a new built-in tactic predict_steps_with_suggestion within the standard Lean user interface. It is a helpful tool that not only suggest tactics but also offer users immediate feedback on proof progress and potential next steps.
@@ 2 Related Work
LLMs for Formal Proof Generation. Large Language Models (LLMs) have demonstrated significant potential in the field of formal theorem proving (Yang et al., 2024a), finding applications across various proof assistants (Yang et al., 2024b; Song et al., 2024; Lama et al., 2024). Current research on LLM-based theorem proving primarily focuses on several key tasks. A prominent application is tactic suggestion. Following GPT-f (Polu & Sutskever, 2020), LLMs are employed to predict the most promising next tactic given the current proof state. These methods are often coupled with proof search algorithms, such as best-first search (Yang et al., 2024b) or majority voting (Zhou et al., 2024), to explore the proof space and discover complete proofs (Wu et al., 2024). Other techniques, such as retrieval-augmented LLMs (Yang et al., 2024b) and agentic approaches (Thakur et al., 2024), provides further aids for tactic generation by selecting relevant lemmas and enabling multi-round proof refinement utilizing environment feedback. Moreover, emerging research directions include autoformalization (Wu et al., 2022; Jiang et al., 2023), which aims to translate informal mathematical text into formal proofs, and the direct generation of complete proof sketches (Jiang et al., 2022; Wang et al., 2024), both of which can be combined with proof generation to enable large-scale training despite inherent proof data scarcity. Our work addresses the gap from local tactic prediction to a global understanding of the proof trajectory by focusing on predicting the number of remaining steps required for proof completion, offering a novel way for new applications of reinforcement learning in automated theorem proving.
@@
Interactive Tools for Formal Theorem Proving. Mathematicians proving theorems in Lean can significantly benefit from interactive tools that integrate seamlessly into the Lean workflow and provide aids. LLMStep (Welleck & Saha, 2023) extracts current proof states from Lean and sends it to a remote server for LLM-generated tactic suggestions. LeanCopilot (Song et al., 2024) improves the user experience by having fully native tactic suggestion and proof search tools in Lean, besides an additional functionality of premise selection, providing more comprehensive assistance for the proving process. CoqPilot (Kozyrev et al., 2024), a VS Code extension for Coq, uses LLMs, among other generative methods, to fill in proof holes by an “admit” tactic. Unlike these tactics or proof-centric approaches, we predict the number of remaining steps by adding a new tactic, predict_steps_with_suggestion based on LeanCopilot, providing tactic suggestions ranked by the output number of remaining steps as a score.
@@
LLM Guidance in Search. Effective proof search is essential for automated theorem proving. While scaling computational resources during search has led to significant advancements, as seen in AlphaGeometry (Trinh et al., 2024) and AlphaProof (AlphaProof & AlphaGeometry, 2024) for IMO problems and in recent work on natural language reasoning (Lightman et al., 2023; Yang et al., 2022; Zhang et al., 2024; Xie et al., 2024) (including OpenAI’s o1, o3 model (Jaech et al., 2024; Xu et al., 2025)), proof search is a bit different. The vast search space of possible proof steps necessitates effective guiding mechanisms. This highlights the need for methods that can provide a global perspective on proof progress, which our work addresses by predicting the number of remaining steps.
@@ 3 Data Generation for LeanProgress
This section details the data generation and processing methodology used to train LeanProgress. We describe the process of generating proof trees using best-first search (BFS) and the Reprover model, the resulting dataset of proof trajectories, and the adjustments made to address the skewed distribution of proof lengths.
@@ 3.1 Preliminaries: Tactic Prediction as an MDP
The theorem-proving problem can be formalized as a Markov Decision Process (MDP), denoted as $(S,A,P_{a},R_{a})$, where $S$ represents the set of all possible proof states. $A$ represents the set of all available tactics (actions). $P_{a}$ represents the state transition probabilities after executing tactic $a$ in state $s$. $R_{a}$ represents the reward obtained by executing tactic $a$. From an MDP perspective, a proof process can be viewed as a trajectory of states, tactics, and rewards: $(s_{i},a_{i},r_{i})$, where the proof assistant (e.g., Lean) provides the next state $s_{i+1}$ given the current state $s_{i}$ and the applied tactic $a_{i}$.
In typical tactic prediction, proving a theorem involves providing a proof state $s$ to a language model $L$, which then generates a tactic $a$, i.e., $\pi_{L}(a|s)$. Typically, final states (where the goal is proven) are assigned a reward of 1, indicating successful completion of the proof.
@@ 3.2 Generating Proof Trees and Trajectories
A common evaluation strategy for neural theorem provers is best-first search (BFS), as used in GPT-f and related research (Han et al., 2021). This method explores the proof space by iteratively expanding the ”best” state, determined by the maximum cumulative log probability of the preceding proof trajectory. Specifically, given a set of unexpanded states $s_{i}$, the ”best” state to expand is chosen according to: $\max_{i}\sum_{j=0}^{i-1}\log p(a_{j},s_{j}),$ where $(s_{0},a_{0}),\dots,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average log probability of the generated tokens for the tactic $a_{j}$ in state $s_{j}$.
Our work utilizes BFS in conjunction with the Reprover model (Yang et al., 2024b) to generate successful proof trees. By systematically applying Reprover to all reachable states within a certain depth in a best-first manner, we construct a tree of successful proofs. This approach allows us to collect a dataset of complete proof trajectories, which is then used to train our model to predict the number of remaining steps. This data generation process is crucial for training our model to understand the relationship between proof states and the number of steps required for completion. In particular, if multiple proofs are found for a theorem (i.e., multiple no_goals nodes are reached), we select the proof with the minimum depth (the length of the path from the root node to the no_goals node) to ensure the quality and consistency of the training data.
Formally, let $T=\{t_{1},t_{2},...,t_{M}\}$ be the set of theorems in our dataset. For each theorem $t_{i}\in T$, we perform BFS using the Reprover model to generate a set of successful proof trajectories $P_{i}=\{p_{i,1},p_{i,2},...,p_{i,k_{i}}\}$, where $k_{i}$ is the number of proofs found for theorem $t_{i}$. Each proof trajectory $p_{i,j}$ is a sequence of proof states: $p_{i,j}=(s_{i,j,1},s_{i,j,2},...,s_{i,j,n_{i,j}})$, where $n_{i,j}$ is the length (number of steps) of the $j$-th proof for theorem $t_{i}$.
If $k_{i}>1$, we select the proof trajectory $p_{i,j^{*}}$ with the minimum depth:
$j^{*}=\arg\min_{j}\{n_{i,j}\mid 1\leq j\leq k_{i}\}$
@@ 3.3 Data Balancing
We evaluated our models on a dataset of Lean proofs extracted from Lean Workbook Plus and Mathlib4. The original dataset exhibited a skewed distribution of proof lengths, with an average proof length of $L_{\text{original}}=2.47$. This distribution is shown in Figure LABEL:fig:original_proof_length_distribution. We adjusted the data distribution based on relative progress within each proof to address this imbalance and ensure a more representative sample of different proof stages.
The adjustment was performed by assigning different sampling ratios to five ranges of proof lengths: 1-5 steps (Basic progress, 0.01 ratio), 6-10 steps (Intermediate progress, 0.3 ratio), 11-15 steps (Moderate progress, 0.5 ratio), 16-20 steps (Advanced progress, 0.7 ratio), and 21+ steps (Expert progress, 1.0 ratio). This strategy effectively upsamples longer proofs and downsamples shorter ones, resulting in a more balanced dataset. The resulting adjusted distribution has an average proof length of $L_{\text{adjusted}}=10.1$, and the comparison is shown in Figure LABEL:fig:adjusted_proof_length_distribution.
The dataset was then split into training, validation, and test sets. The test set contains 88,233 proof states. The training set contains $N_{\text{train}}$ proof states, and the validation set contains $N_{\text{val}}$ proof states. The dataset was partitioned randomly at the theorem level, meaning that all states from a given theorem belong to the same split (either training, validation, or test). This prevents data leakage between splits.

### v3

@@ Abstract
Mathematical reasoning remains a significant challenge for Large Language Models (LLMs) due to hallucinations. When combined with formal proof assistants like Lean, these hallucinations can be eliminated through rigorous verification, making theorem proving reliable. However, even with formal verification, LLMs still struggle with long proofs and complex mathematical formalizations. While Lean with LLMs offers valuable assistance with retrieving lemmas, generating tactics, or even complete proofs, it lacks a crucial capability: providing a sense of proof progress. This limitation particularly impacts the overall development efficiency in large formalization projects. We introduce LeanProgress, a method that predicts the progress in the proof. Training and evaluating our models made on a large corpus of Lean proofs from Lean Workbook Plus and Mathlib4 and how many steps remain to complete it, we employ data preprocessing and balancing techniques to handle the skewed distribution of proof lengths. Our experiments show that LeanProgress achieves an overall prediction accuracy of 75.8% in predicting the amount of progress and, hence, the remaining number of steps. When integrated into a best-first search framework using Reprover, our method shows a 3.8% improvement on Mathlib4 compared to baseline performances of 41.4%, particularly for longer proofs. These results demonstrate how proof progress prediction can enhance both automated and interactive theorem proving, enabling users to make more informed decisions about proof strategies. Our code is merged in this library here https://github.com/lean-dojo/LeanDojo-v2.
@@ 1 Introduction
Formal theorem proving (Avigad, 2023) has emerged as a cornerstone of rigorous mathematical verification, providing machine-checked guarantees for proofs ranging from foundational results (Gowers et al., 2023) to industrial applications (Community, 2022). The Lean proof assistant (Moura and Ullrich, 2021), built on dependent type theory, has witnessed remarkable adoption growth (Best et al., 2023), fueled by collaborative efforts on large-scale formalization projects (mathlib Community, 2020) and novel mathematical developments (Asgeirsson, 2024). This collaborative paradigm shift underscores the urgent need for enhanced tooling to support mathematicians navigating increasingly complex proof environments.
The recent success of Large Language Models (LLMs) in code generation (Rozière et al., 2024) and symbolic reasoning (Yu et al., 2024) has spurred innovations at the intersection of LLMs and formal verification. Some of the works that have been developed include LeanDojo (Yang et al., 2024b) which provides an interactive environment for training LLMs on tactic-level interactions while LLMStep (Welleck and Saha, 2023) and LeanCopilot (Song et al., 2024) focuses on next-tactic suggestion through interface as a useful tool. Lean Agent (Kumarappan et al., 2024) then combines neural suggestion with life-long learning while Lean Aide (Agrawal et al., 2022) and Lean-STaR(Lin et al., 2024) translates statements written in natural language in a doc-string like format to Lean types (including theorem statements) and bootstrapping thoughts. While these systems demonstrate impressive tactic-level accuracy (Johansson, 2023), they primarily optimize for local correctness rather than global proof progress – a critical limitation when navigating Lean’s vast action space (Nawrocki et al., 2023).
@@
Reinforcement learning (RL) presents a theoretically appealing framework for automated theorem proving (Dong et al., 2024), where finding reward signals over proof trajectories is essential. However, the combinatorial explosion of tactic sequences in Lean (Clune, 2023) renders direct RL applications impractical (Setlur et al., 2024). Alphaproof (Hubert et al., 2025) has done RL for theorem proving but it’s not open source and needs enormous compute. Current approaches mitigate this through hybrid architectures (Wang et al., 2023) but remain fundamentally limited by the absence of reliable progress indicators to guide exploration – a prerequisite for effective RL in mathematical domains (Gao et al., 2024).
@@
LeanProgress makes the following key contributions: Balanced Data Generation Pipeline: We construct a balanced dataset of approximately 80k proof trajectories from Lean Workbook Plus and Mathlib4 by performing a tree search and selecting the shortest path as ground truth. We employ a data balancing strategy based on relative proof progress. Since the useful and non-trivial data of long proofs are long-tailed distributed in the original dataset, we fully utilize long proof data by assigning each state with a remaining step as a label.
Model for Progress Prediction: We fine-tune a DeepSeek Coder V1 1.3b base model to predict the remaining steps, achieving a Mean Absolute Error (MAE) of 3.15 and an overall prediction accuracy of 75.8% on the test set with proof history. Unlike tactic suggestion tools, LeanProgress provides a global view of the proof process by predicting the remaining steps rather than the immediate next tactic.
Progress-Guided Proof Search: We integrate our step prediction model into a best-first search framework. A natural first step for using the progress predictor is combining the predicted remaining steps with the tactic generator’s log probabilities to guide the search. In the future, we hope to use this, instead of just relying on the log probabilities, as a reward for RL. We observe on Mathlib4 a significant improvement of 3.8% with the baseline Reprover performance of 41.4%.
Integration with LeanCopilot: Based on the LeanCopilot framework, we provide a new built-in tactic predict_steps_with_suggestion within the standard Lean user interface. It is a helpful tool that not only suggest tactics but also offer users immediate feedback on proof progress and potential next steps. All the development details are in Appendix B and C.
@@ 2 Related Work
LLMs for Formal Proof Generation. Large Language Models (LLMs) have demonstrated significant potential in the field of formal theorem proving (Yang et al., 2024a), finding applications across various proof assistants (Yang et al., 2024b; Song et al., 2024; Lama et al., 2024). Current research on LLM-based theorem proving primarily focuses on several key tasks. A prominent application is tactic suggestion. Following GPT-f (Polu and Sutskever, 2020), LLMs are employed to predict the most promising next tactic given the current proof state. These methods are often coupled with proof search algorithms, such as best-first search (Yang et al., 2024b) or majority voting (Zhou et al., 2024), to explore the proof space and discover complete proofs (Wu et al., 2024). Other techniques, such as retrieval-augmented LLMs (Yang et al., 2024b) and agentic approaches (Thakur et al., 2024), provides further aids for tactic generation by selecting relevant lemmas and enabling multi-round proof refinement utilizing environment feedback. Moreover, emerging research directions include autoformalization (Wu et al., 2022; Jiang et al., 2023), which aims to translate informal mathematical text into formal proofs, and the direct generation of complete proof sketches (Jiang et al., 2022; Wang et al., 2024), both of which can be combined with proof generation to enable large-scale training despite inherent proof data scarcity. Our work addresses the gap from local tactic prediction to a global understanding of the proof trajectory by focusing on predicting the number of remaining steps required for proof completion, offering a novel way for new applications of reinforcement learning in automated theorem proving.
@@
Interactive Tools for Formal Theorem Proving. Mathematicians proving theorems in Lean can significantly benefit from interactive tools that integrate seamlessly into the Lean workflow and provide aids. LLMStep (Welleck and Saha, 2023) extracts current proof states from Lean and sends it to a remote server for LLM-generated tactic suggestions. LeanCopilot (Song et al., 2024) improves the user experience by having fully native tactic suggestion and proof search tools in Lean, besides an additional functionality of premise selection, providing more comprehensive assistance for the proving process. CoqPilot (Kozyrev et al., 2024), a VS Code extension for Coq, uses LLMs, among other generative methods, to fill in proof holes by an “admit” tactic. Unlike these tactics or proof-centric approaches, we predict the number of remaining steps by adding a new tactic, predict_steps_with_suggestion based on LeanCopilot, providing tactic suggestions ranked by the output number of remaining steps as a score.
@@
LLM Guidance in Search. Effective proof search is essential for automated theorem proving. While scaling computational resources during search has led to significant advancements, as seen in AlphaGeometry (Trinh et al., 2024) and AlphaProof (Hubert et al., 2025) for IMO problems and in recent work on natural language reasoning (Lightman et al., 2023; Yang et al., 2022; Zhang et al., 2024; Xie et al., 2024) (including OpenAI’s o1, o3 model (Jaech et al., 2024; Xu et al., 2025)), proof search is a bit different. The vast search space of possible proof steps necessitates effective guiding mechanisms. This highlights the need for methods that can provide a global perspective on proof progress, which our work addresses by predicting the number of remaining steps.
Progress Signals for Proof Search. Several systems learn value-like guidance signals for proof search. GamePad (Huang et al., 2019) introduces a learning environment for Coq and studies both tactic prediction and position evaluation, producing estimates of how promising a proof state is. More recently, QEDCartographer (Sanchez-Stern et al., 2025) uses reinforcement learning to learn a proof-state evaluation signal tailored to the branching structure of proofs, improving automated proof synthesis in Coq. LeanProgress targets a complementary notion of guidance: a supervised, interpretable distance-to-goal signal by predicting the remaining number of steps to completion from successful traces. Unlike generic value estimates, this output is directly interpretable as progress, and it composes naturally with existing tactic generators (e.g., as a lightweight reranking prior alongside log-likelihood).
@@ 3 Data Generation for LeanProgress
This section details the data generation and processing methodology used to train LeanProgress. We describe the process of generating proof trees using best-first search and the Reprover model, the resulting dataset of proof trajectories, and the adjustments made to address the skewed distribution of proof lengths.
@@ 3.1 Preliminaries: Tactic Prediction as an MDP
The theorem-proving problem can be formalized as a Markov Decision Process (MDP), denoted as $(S,A,P_{a},R_{a})$, where $S$ represents the set of all possible proof states. $A$ represents the set of all available tactics (actions). $P_{a}$ represents the state transition probabilities after executing tactic $a$ in state $s$. $R_{a}$ represents the reward obtained by executing tactic $a$. From an MDP perspective, a proof process can be viewed as a trajectory of states, tactics, and rewards: $(s_{i},a_{i},r_{i})$, where the proof assistant (e.g., Lean) provides the next state $s_{i+1}$ given the current state $s_{i}$ and the applied tactic $a_{i}$. In typical tactic prediction, proving a theorem involves providing a proof state $s$ to a language model $L$, which then generates a tactic $a$, i.e., $\pi_{L}(a|s)$. Typically, final states (where the goal is proven) are assigned a reward of 1, indicating successful completion of the proof.
@@ 3.2 Generating Proof Trees and Trajectories
A common evaluation strategy for neural theorem provers is best-first search, as used in GPT-f and related research (Han et al., 2021). This method explores the proof space by iteratively expanding the "best" state, determined by the maximum cumulative log probability of the preceding proof trajectory. Specifically, given a set of unexpanded states $s_{i}$, the "best" state to expand is chosen according to: $\max_{i}\sum_{j=0}^{i-1}\log p(a_{j},s_{j}),$ where $(s_{0},a_{0}),\dots,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average log probability of the generated tokens for the tactic $a_{j}$ in state $s_{j}$.
Our work utilizes best-first search in conjunction with the Reprover model (Yang et al., 2024b) to generate successful proof trees. By systematically applying Reprover to all reachable states within a certain depth in a best-first manner, we construct a tree of successful proofs. This approach allows us to collect a dataset of complete proof trajectories, which is then used to train our model to predict the number of remaining steps. This data generation process is crucial for training our model to understand the relationship between proof states and the number of steps required for completion. In particular, if multiple proofs are found for a theorem (i.e., multiple no_goals nodes are reached), we select the proof with the minimum depth (the length of the path from the root node to the no_goals node) to ensure the quality and consistency of the training data.
Formally, let $T=\{t_{1},t_{2},...,t_{M}\}$ be the set of theorems in our dataset. For each theorem $t_{i}\in T$, we perform best-first search using the Reprover model to generate a set of successful proof trajectories $P_{i}=\{p_{i,1},p_{i,2},...,p_{i,k_{i}}\}$, where $k_{i}$ is the number of proofs found for theorem $t_{i}$. Each proof trajectory $p_{i,j}$ is a sequence of proof states: $p_{i,j}=(s_{i,j,1},s_{i,j,2},...,s_{i,j,n_{i,j}})$, where $n_{i,j}$ is the length (number of steps) of the $j$-th proof for theorem $t_{i}$.
If $k_{i}>1$, we select the proof trajectory $p_{i,j^{*}}$ with the minimum depth: $j^{*}=\arg\min_{j}\{n_{i,j}\mid 1\leq j\leq k_{i}\}$
@@
(a) Original dataset of steps distribution
(b) Adjusted dataset of steps distribution by balancing different ranges
Figure 3: Proof‐length distributions before and after length balancing. Panel (a) plots all collected states prior to balancing (dominated by 1–2-step proofs; average length $L_{\text{original}}=2.47$). Panel (b) shows the training set after balancing, where we down-sample short proofs and (lightly) up-sample longer ones, yielding fewer total states and a higher average length $L_{\text{balanced}}=10.1$. Bar heights are raw counts, so the totals differ and the $y$-axes are not directly comparable.
@@ 3.3 Data Balancing
We evaluated our models on a dataset of Lean proofs extracted from Lean Workbook Plus and Mathlib4. The original dataset exhibited a skewed distribution of proof lengths, with an average proof length of $L_{\text{original}}=2.47$. This distribution is shown in Figure 3(a). We adjusted the data distribution based on relative progress within each proof to address this imbalance and ensure a more representative sample of different proof stages. The adjustment was performed by assigning different sampling ratios to five ranges of proof lengths: 1-5 steps (Basic progress, 0.01 ratio), 6-10 steps (Intermediate progress, 0.3 ratio), 11-15 steps (Moderate progress, 0.5 ratio), 16-20 steps (Advanced progress, 0.7 ratio), and 21+ steps (Expert progress, 1.0 ratio). This strategy effectively upsamples longer proofs and downsamples shorter ones, resulting in a more balanced dataset. The resulting adjusted distribution has an average proof length of $L_{\text{adjusted}}=10.1$, and the comparison is shown in Figure 3(b).
The dataset was then split into training, validation, and test sets. The test set contains 8820 proof states. The training set contains $N_{\text{train}}$ proof states, and the validation set contains $N_{\text{val}}$ proof states. The dataset was partitioned randomly at the theorem level, meaning that all states from a given theorem belong to the same split (either training, validation, or test). This prevents data leakage between splits.